OpenStudy (anonymous):

solve -x+y+z=4 4x-5y+2z=0 2x+3y+z=11

OpenStudy (akshay_budhkar):

hi joe lana..

OpenStudy (anonymous):

solve -x+y+z=4 4x-5y+2z=0 2x+3y+z=11

OpenStudy (akshay_budhkar):

i cant help you solving it but the steps and solution is here http://www.wolframalpha.com/input/?i=-x%2By%2Bz%3D4+%2C4x-5y%2B2z%3D0%2C+2x%2B3y%2Bz%3D11+ hope it helps :)

OpenStudy (anonymous):

2x + 3y + z = 11 +2(-x +y +z = 4) => 5y + 3z = 19 4x - 5y + 2z = 0 +4(-x +y +z = 4) => -y + 6z = 16 then it's a simple 2 variables 2 equations which you've been practicing after you solve for y in terms of z or z in terms of y. you use that back into your original equations. and then you use that to start solving for the variables in terms of actual numbers. it takes a lot more manipulation

OpenStudy (lalaly):

z=4+x-y... substitute it in the las two equations 4-5y+8+2x-2y=0>>> 6x-7y=-8...(1) 2x+3y+4+x-y=11>>> 3x+2y=7....(2) now multiply (2) by 2 6x-7y=-8 6x+4y=14 .............................subtract -11y=-22 y=2 6x-7(2)=-8 6x=6 x=1 z=4+x-y = 4+1-2=3

OpenStudy (anonymous):

hashir samjha mujhe ye

OpenStudy (anonymous):

http://www.twiddla.com/589124 yahaan per aajao